OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±6) and asymptotes at y = ±3/2.x.

hartnn (hartnn):

For the standard form of \(\Large \dfrac{y^2 }{a^2 } -\dfrac{x^2 }{b^2 } =1\) the vertices are \(\Large (0, \pm a)\) and you have vertices as (0,\(\pm\)6) so you directly get the value of 'a' as ?

OpenStudy (anonymous):

i'm not sure?

hartnn (hartnn):

Compare \((0,\pm a)\) with \((0,\pm 6)\)

OpenStudy (anonymous):

im sorry, i dont understand :(

hartnn (hartnn):

ok, let me give the steps, see whether you get them, if u have any doubt in any step, then ask :)

OpenStudy (anonymous):

ok

hartnn (hartnn):

\(Vertices \equiv (0,\pm a) \equiv (0,\pm 6) \\ \Large \implies a=6 \) Asymptotes for this hyperbola are \(\Large y = \pm \dfrac{a}{b} x\) and we have \(\Large y = \pm \dfrac{3}{2} x\) so, a/b = 3/2 and we already found a=6 so 6/b =3/2 6*2 =3*b b= 12/3 = 4 so the equation of hyperbola is \(\Large \dfrac{y^2 }{6^2 } -\dfrac{x^2 }{4^2}=1\) or \(\Large \dfrac{y^2 }{36 } -\dfrac{x^2 }{16}=1\) thats it! :)

OpenStudy (anonymous):

O.o that looks simple enough with the formulas. thanks

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

do you think you could help me with a few more?